A man has extra digits (six fingers on each hand and six toes on each foot). His wife and their daughter have a normal number of digits. Having extra digits is a dominant trait. The couples second child has extra digits. What is the probability that their next (third) child will have extra digits? 1. 1/8 2. 1/2 4.9/16 5.3/4
Answer #1

Man = Dd

Woman = dd

Dd × dd

Gametes d
D Dd extra fingers
d dd normal fingers

Probability of extra fingers = 1/2

Probability of normal fingers = 1/2

First child is normal, second child has extra digits and third child is also havin extra digits.

Probability of third child to have extra digits = 1/2 × 1/2 × 1/2 = 1/8

Answer is 1/8.

Please rate high.